# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ A Pauli error channel is given by
We can compare the $\langle Z_{\textrm{tot}}^{2}\rangle$ values to the exact state vector simulation results for small system sizes for all steps and to the H2 data with ZNE + ZNR for 7x8 system. The results with the largest bond dimension $D=10$ start to deviate notably from Trotter step 5 onwards for all system sizes. The deviation has a similar trend as MPS with respect to overestimating the $\langle Z_{\textrm{tot}}^{2}\rangle$ value. Separating the system into two by a horizontal cut at $L_{y}/2$, the half-cut Rényi-2 entanglement entropy $S^{[2]}$ can be estimated with the reduced density matrix formed by the boundary MPO. For states with 2D area-law entanglement, the rescaled entanglement entropies $S^{[2]}/L_{x}$ should remain constant with increasing system sizes. In Fig. A17, we observe a similar convergence trend of entanglement entropies compared to $\langle Z_{\textrm{tot}}^{2}\rangle$, where data converged nicely until Trotter step 3. The entanglement entropies with the largest bond dimension $D=10$ also deviate notably from Trotter step 5 onwards for the 4x4 system. The maximal values of the rescaled entanglement entropies for system sizes 6x6 and 7x8 exceed the maximal value for system size 4x4, indicating the entanglements are beyond area law.
@@
One potential way forward is to utilize the ansatz in an unconventional way that goes to larger bond dimensions by performing approximate updates, e.g., SU. In the next section, we discuss the Belief Propagation update approach, which is a class of methods generalizing the SU approach. Another effect we have to consider is that using an OBC-PEPS to simulate periodic boundary conditions (PBC) dynamics itself has a significant overhead in the bond dimension, as the correlations have to be carried from one of the open ends to the other. A heuristic argument based on collapsing the PBC-PEPS to an OBC-PEPS suggests that it requires a bond dimension $D^{2}$ OBC-PEPS to represent the same state represented by a bond dimension $D$ PBC-PEPS. Again, we discuss this issue and consider this effect of the ansatz in the coming section.
@@
Using BP, we can efficiently obtain the time-evolved PEPS with a considerably larger bond dimension $D$ than is possible with the FU method. However, the main challenge lies in measuring the observable $\langle Z_{\textrm{tot}}^{2}\rangle$, which remains the most difficult task in the process. While several strategies exist, we explored the loop series expansion [80] and clustering approximation [81]. However, we ultimately achieved the best results using Monte Carlo sampling. The latter procedure works as follows: for a given Trotter step, we obtain an approximate MPS wavefunction with bond dimension $D=520$ (as described in Sec. A4), and then efficiently produce samples $|x\rangle$ with probability $w_{\text{MPS}}(x)$ using perfect sampling [82]. We then estimate the true probability $w_{\rm PEPS}(x)=|\bra{x}\Psi_{\rm PEPS}\rangle|^{2}$ by approximately contracting the associated tensor network, and use such probabilities to form a Monte Carlo estimate of observables. This method provides an unbiased estimate of an observable $\bra{\Psi_{\rm PEPS}}O\ket{\Psi_{\rm PEPS}}$ (assuming no approximation error in computing $|\bra{x}\Psi_{\rm PEPS}\rangle|^{2}$) even if the MPS forms a poor approximation to |$\ket{\Psi_{\rm PEPS}}$, but the quality of the MPS facilitates a degree of importance sampling (and therefore improves estimate variances). Note that we expect $w_{\text{MPS}}(x)$ to serve as a reasonably good approximation of the true PEPS probability distribution $w_{\text{PEPS}}(x)$, thereby reducing sample overhead (though this breaks down at late times, see Fig. A18). Measuring the observable using importance sampling is significantly simpler than direct estimation (requiring double-layer PBC-PEPS approximate contraction), as it requires only a single-layer PBC-PEPS, as shown here,
[the notation $x\sim w_{\rm MPS}(x)$ in the sum subscript implying that the sum samples $x$ from the distribution $w_{\rm MPS}(x)$]. The single-layer PBC-PEPS shown [the square of which represents $w_{\rm PEPS}(x)$] is estimated using a hyper-optimized approximate contraction approach, controlled by a boundary bond dimension $\chi$ [83]. Empirically, it is found that a boundary bond dimension $\chi=2D$ is sufficient to produce converged results, comparable to the sampling overhead (see Fig. A19a).
@@
Recent works [40, 89] have demonstrated the viability of neural-network quantum state (NQS) [90], when combined with time-dependent variational Monte Carlo (tVMC) [41], as a competitive tool for simulating quantum many-body dynamics compared to the state of the art infinite PEPS simulation [91, 92]. These works have served as concrete benchmarks and stimulated the development of new algorithms in simulating dynamics with NQS, for example, the projected time-dependent variational Monte Carlo (ptVMC) method [93, 94, 95, 96, 97, 98] and the time-dependent neural quantum state (t-NQS) [99] method.
@@
We first present the results for the $5\times 5$ system with periodic boundary conditions, alongside the exact results, in Fig. A24, panels (a) and (b). For the single-layer CNN, we observe that increasing the number of features reduces the error in the observables. The two-layer CNN requires significantly longer computational times and yields less accurate results. Consequently, we focus on the single-layer architecture for larger system sizes. We also observe that single-site observables are easier to capture accurately than two-point correlation functions. This is particularly evident in panels (a) and (b) of Fig. A24. The simulation with (4,3) features for the 5x5 system yields reasonably accurate results for the magnetization up to time step 8, whereas the values obtained for the magnetization squared show significant deviations as early as time step 4.

### v2

@@ A Pauli error channel is given by
We can compare the $\langle Z_{\textrm{tot}}^{2}\rangle$ values to the exact state vector simulation results for small system sizes for all steps and to the H2 data with ZNE + ZNR for $7\times 8$ system. The results with the largest bond dimension $D=10$ start to deviate notably from Trotter step 5 onwards for all system sizes. The deviation has a similar trend as MPS with respect to overestimating the $\langle Z_{\textrm{tot}}^{2}\rangle$ value. Separating the system into two by a horizontal cut at $L_{y}/2$, the half-cut Rényi-2 entanglement entropy $S^{[2]}$ can be estimated with the reduced density matrix formed by the boundary MPO. For states with 2D area-law entanglement, the rescaled entanglement entropies $S^{[2]}/L_{x}$ should remain constant with increasing system sizes. In Fig. A17, we observe a similar convergence trend of entanglement entropies compared to $\langle Z_{\textrm{tot}}^{2}\rangle$, where data converged nicely until Trotter step 3. The entanglement entropies with the largest bond dimension $D=10$ also deviate notably from Trotter step 5 onwards for the $4\times 4$ system. The maximal values of the rescaled entanglement entropies for system sizes $6\times 6$ and $7\times 8$ exceed the maximal value for system size $4\times 4$, indicating the entanglements are beyond area law.
@@
One potential way forward is to utilize the ansatz in an unconventional way that goes to larger bond dimensions by performing approximate updates, e.g., SU. In the next section, we discuss the Belief Propagation update approach, which is a class of methods generalizing the SU approach. Another effect we have to consider is that using an OBC-PEPS to simulate periodic boundary conditions (PBC) dynamics itself has a significant overhead in the bond dimension, as the correlations have to be carried from one of the open ends to the other. A heuristic argument based on collapsing the PBC-PEPS to an OBC-PEPS suggests that it requires a bond dimension $D^{2}$ OBC-PEPS to represent the same state corresponding to a bond dimension $D$ PBC-PEPS. Again, we discuss this issue and consider this effect of the ansatz in the coming section.
@@
Using BP, we can efficiently obtain the time-evolved PEPS with a considerably larger bond dimension $D$ than is possible with the FU method. However, the main challenge lies in measuring the observable $\langle Z_{\textrm{tot}}^{2}\rangle$, which remains the most difficult task in the process. While several strategies exist, we explored the loop series expansion [82] and clustering approximation [83]. However, we ultimately achieved the best results using Monte Carlo sampling. The latter procedure works as follows: for a given Trotter step, we obtain an approximate MPS wavefunction with bond dimension $D=520$ (as described in Sec. A4), and then efficiently produce samples $|x\rangle$ (in the computational basis) with probability $w_{\text{MPS}}(x)$ using perfect sampling [84]. We then estimate the true probability $w_{\rm PEPS}(x)=|\bra{x}\Psi_{\rm PEPS}\rangle|^{2}$ by approximately contracting the associated tensor network, and use such probabilities to form a Monte Carlo estimate of observables. This method provides an unbiased estimate of an observable $\bra{\Psi_{\rm PEPS}}O\ket{\Psi_{\rm PEPS}}$ (assuming no approximation error in computing $|\bra{x}\Psi_{\rm PEPS}\rangle|^{2}$) even if the MPS forms a poor approximation to |$\ket{\Psi_{\rm PEPS}}$, but the quality of the MPS facilitates a degree of importance sampling (and therefore improves estimate variances). Note that we expect $w_{\text{MPS}}(x)$ to serve as a reasonably good approximation of the true PEPS probability distribution $w_{\text{PEPS}}(x)$ at short times, thereby reducing sample overhead due to effective importance sampling. Figure A18 quantifies the quality of the distribution $w_{\rm MPS}(x)$ with respect to $w_{\rm PEPS}(x)$ via the Kullback-Leibler (KL) divergence $D_{\rm KL}(w_{\rm MPS}||w_{\rm PEPS})$. Small values of $D_{\rm KL}$ at small $s$ indicate good overlap between $w_{\rm MPS}$ and $w_{\rm PEPS}$, implying effective importance sampling. Larger values of $D_{\rm KL}$ at larger $s$ indicate the breakdown of importance sampling, leading to larger estimate variances.
Measuring the observable using importance sampling is significantly simpler than direct estimation (requiring double-layer PBC-PEPS approximate contraction), as it only requires the contraction of a single-layer PBC-PEPS,
Here, $Z_{x}^{2}=\langle x|Z_{\textrm{tot}}^{2}|x\rangle$ is simple to compute as the samples |$\ket{x}$ are in the computational basis (i.e., each |$\ket{x}$ is an eigenstate of all $Z_{j}$). The notation $x\sim w_{\rm MPS}(x)$ in the sum subscript implies that the sum samples $x$ from the distribution $w_{\rm MPS}(x)$, and the single-layer PBC-PEPS diagram in Eq. (A21) represents the amplitude $\langle x\ket{\Psi_{\rm PEPS}}$ [its square represents $w_{\rm PEPS}(x)$]. We estimate this diagram using a hyper-optimized approximate contraction approach, controlled by a boundary bond dimension $\chi$ [85]. Empirically, it is found that a boundary bond dimension $\chi=2D$ is sufficient to produce converged results, comparable to the sampling overhead (see Fig. A19a).
@@
Recent works [40, 91] have demonstrated the viability of neural-network quantum state (NQS) [92], when combined with time-dependent variational Monte Carlo (tVMC) [41], as a competitive tool for simulating quantum many-body dynamics compared to the state of the art infinite PEPS simulation [93, 94]. These works have served as concrete benchmarks and stimulated the development of new algorithms in simulating dynamics with NQS, for example, the projected time-dependent variational Monte Carlo (ptVMC) method [95, 96, 97, 98, 99, 100] and the time-dependent neural quantum state (t-NQS) [101, 102] method.
@@
We first present the results for the $5\times 5$ system with periodic boundary conditions, alongside the exact results, in Fig. A24, panels (a) and (b). For the single-layer CNN, we observe that increasing the number of features reduces the error in the observables. The two-layer CNN requires significantly longer computational times and yields less accurate results. Consequently, we focus on the single-layer architecture for larger system sizes. We also observe that single-site observables are easier to capture accurately than two-point correlation functions. This is particularly evident in panels (a) and (b) of Fig. A24. The simulation with (4,3) features for the $5\times 5$ system yields reasonably accurate results for the magnetization up to time step 8, whereas the values obtained for the magnetization squared show significant deviations as early as time step 4.